# Structured Packing in LLM Training Improves Long Context Utilization

### v2

@@ 4 Large-scale models
In this section, we show that SPLiCe can improve the long context performance of large-scale language models. To this end, we use $3$B parameter models and tasks that test in-context learning, question answering, and information retrieval capabilities. We provide perplexity results in Appendix E
@@ 4.1 Experimental setup
We continue training OpenLLaMA 3B v2 (Geng & Liu, 2023; Geng, 2023) on the dataset defined in Section 3.1 - that is, a 50/50 mixture of RedPajama prepared in a standard way and C prepared using a method of choice. Specifically, we train two models with $32K$ context length for the same number of tokens, 5.4B, which is $21$k training steps, on two datasets: one generated using SPLiCe, and the other one using the Baseline, see Section 2 for definitions. We use a batch size of 256K tokens per step, and learning rate of ${1.5}\mathrm{e}{-5}$ with linear warmup and cosine decay, following (Geng & Liu, 2023). We evaluate the perplexity on a mixture of long documents sourced from CommonCrawl, arXiv, and StarCoder (Li et al., 2023b). Otherwise, we follow the protocol described in Section 3.1.
4.1.1 In context learning
To evaluate the improvements of the model’s in-context learning ability we use TREC (Li & Roth, 2002; Hovy et al., 2001) question classification task. We provide the SPLiCe trained model with $190$ to $1560$ in-context examples and compare the prediction accuracy against the Baseline model. We sample the few-shot examples using 50 different random seeds and show the histogram of accuracy improvements in Figure 3 and additional results in Table 4 and Figure 7.
Figure 3: Histogram of accuracy improvement of SPLiCe BM25 over Baseline on TREC question classification task. The results are obtained by comparing the accuracy on the test set of TREC of the 3B model trained with SPLiCe to the model trained with default data preparation method (Baseline) across $50$ sets of in-context examples. Each set of in-context examples consists of $1560$ elements (roughly $32$k tokens) randomly sampled (without replacement) from the training subset of TREC.
4.2 Standardized Comparison Over Long Language Sequences
We show that SPLiCe can improve model performance on Qasper (Dasigi et al., 2021; Shaham et al., 2022). Qasper tests the model’s ability to answer questions regarding research papers provided as a part of the input. The evaluated models were trained on a mixture of RedPajama prepared in a standard way and C prepared either in a standard way (Baseline) or using SPLiCe. Note that despite SPLiCe using only code data here, we still see improvements in score on non-code, see Table 5.
@@ 4.2 Standardized Comparison Over Long Language Sequences
Table 5: Zero-shot performance on the validation subset of Qasper (Dasigi et al., 2021) and few-shot performance on TREC. For Qasper we use the implementation from Language Model Evaluation Harness (Gao et al., 2021). In the Harness implementation, yes/no questions are evaluated separately from open questions. Note that despite using SPLiCe for code data only we still have improvements in non-code tasks. For TREC we average results across $50$ sets of in-context examples. Note that for TREC SPLiCe model is better for almost all choices of in-context examples (see Figures 3 and 7).
Task Context length 3B Baseline 3B SPLiCe
TREC 32K 73.9 $\pm$3.9 79.3 $\pm$ 2.9
16K 68.9 $\pm$5.9 76.9 $\pm$ 3.1
8K 66.5 $\pm$6.2 75.8 $\pm$ 3.5
4K 65.1 $\pm$6.1 72.8 $\pm$ 4.9
Qasper F1 32K 23.1 23.9
4K 18.6 18.7
4.3 Key Retrieval Performance
@@
Recent works indicated numerous issues related to long-context utilization. In particular, (Liu et al., 2023, Figure 7) proposed a benchmark retrieval task to measure the ability of language models to retrieve information from a long context. Specifically, the model is presented with a list of key-value pairs, which are $128$-bit UUIDs, and is tasked to retrieve the value for a given key. Even though very simple, the task proved to be challenging for many open-source language models.
In Figure 4 we present the key retrieval performance of the models trained with SPLiCe and Baseline. The dictionary contains $75$ pairs, which is about 6k tokens. The performance depends heavily on the key position (which is also the case for other models, as indicated in (Liu et al., 2023)). The keys near the beginning are much harder to retrieve, and SPLiCe improves the accuracy more than two times.
@@
Figure 4: Key-value retrieval performance on a dictionary of $75$ key-value pairs ($\approx$ 6k tokens). The model trained with SPLiCe achieves much higher accuracy on hard-to-retrieve positions. Both models achieve $100$% accuracy near the end because the desired key is in the local context (below $2048$ tokens). The details about this task can be found in Appendix D.
@@ 5 Related work
Some recent studies have focused on enhancing training examples to improve the performance of language models. For instance, (Levine et al., 2022) developed a theory and demonstrated empirically that incorporating non-adjacent but semantically related sentences in training examples leads to better sentence embeddings and improves open-domain question-answering performance. Another work (Gu et al., 2023) introduced a pretraining framework grounded on the idea that text documents often include ”intrinsic tasks”. They showed that this approach substantially boosts in-context learning. Additionally, there is existing work on training long-context language models using repository-level code data, such as (Wu et al., 2022).
Our methodology diverges from these works in several key ways. Firstly, we focus on the document-level context during the training phase, as opposed to sentence-level (Levine et al., 2022) or paragraph-level (Gu et al., 2023) granularity. We demonstrate the efficacy of this approach in large-scale language modeling, specifically with OpenLLaMA 3B. Secondly, we construct a tree structure of related documents through BM25/Contriever-MSMARCO retrieval and linearize this structure to form long-context examples. This allows for greater control over example coherence compared to relying solely on natural data structures like repository-level code.
@@ 6 Limitations and future work
We show that structuring the training data is a viable way of improving the model’s performance. The presented method can viewed as a general framework for organizing the documents into training examples (i.e. to achieve the repository-level grouping, it suffices to replace the $\texttt{RETRIEVE}(d,k)$ method with one that returns files from the closest catalog in the repository). This opens multiple further research avenues.
@@
In this work, SPLiCe constructed training examples by concatenating most matching documents (according to BM25/Contriever-MSMARCO). However, recent results of (Tworkowski et al., 2023) show that introducing unrelated data to the context may help the model learn better representations. We leave the study of how the choice of retriever (in particular, how the ratio of related and unrelated documents) affects performance for future work. Note that as BM25 is a text-statistic based retriever, it is not guaranteed that the documents returned by it are semantically related. Moreover, we have not tuned Contriever-MSMARCO, and it was provided only with a $512$ token long prefix of a document.
Another, avenue for future work is to study the granularity of the pieces from which the training examples are constructed. In this work, we focus on the document-level granularity. However, it is possible to construct training examples from smaller pieces, such as paragraphs or sentences.
Our method requires further studies to understand its scaling properties beyond the range presented in the paper. In particular, we leave scaling the context length to hundreds of thousands of tokens and experiments with models having more than $3$B parameters for future work.
Our models were trained for a long context using the methodology presented in (Tworkowski et al., 2023). This methodology is similar to the approach used in (Wu et al., 2022) and allows extension of the model context with relatively small (compared to other methods) computational cost. Briefly speaking, it boils down to choosing a subset of attention layers for the context extension. We have additionally tested three popular context extension methods on a medium-scale (Naive, YaRN, and CodeLlama). We leave studies of SPLiCe with other context extension methods as future work.
One of the approaches to training long-context language models is to use long conversational data (Li et al., 2023a). This approach is complementary to our method. SPLiCe can utilize data that already exists in vast quantities and can be easily applied to different types of text (like code, Wikipedia articles, StackExchange questions, and answers, etc.) to further increase the number of long-context examples. We leave researching how SPLiCe integrates with other methods for preparing the long-context data as future work.
@@
In order to prepare the training data SPLiCe uses each document exactly once (we mask out each used document for further retrieval). However, it is possible that allowing some high-quality documents to occur more than once may be beneficial.
@@
Using highly correlated samples has the potential to result in training instability. However, during our experiments with StarCoder, Wikipedia, and StackExchange data we have noted no performance degradation. We leave the study of how SPLiCe integrates with different types of data for the future. In particular, in our studies used datasets were reasonably deduplicated.
@@ 7 Conclusions
In this work, we present SPLiCe, a novel method of constructing training examples for long-context language models, which utilizes BM25/Contriever-MSMARCO to find relevant documents and feed them to the model in a structured manner. We show that SPLiCe improves the perplexity of on language modeling task and improves the performance on downstream tasks. We further show that SPLiCe can be used to train large-scale models, resulting in a model with improved long-context utilization. Besides its direct usefulness, we believe that our work indicates multiple interesting research directions on how to improve the performance of long-context language models with structured data.
8 Reproducibility
To ensure the reproducibility of our results we provide the details about how the data was prepared in Appendix F, the pseudocode of our method in Algorithm 1 and details about the trained models in Appendix A. We also attach the source code in the supplementary material.
@@ Appendix A Architecture
The architecture of our models is based on LLaMA (Touvron et al., 2023), and the architectural details can be found in Table 6. Briefly speaking, our architecture is similar to the one introduced in (Vaswani et al., 2017) with a few standard changes. First, we use only the decoder without the encoder part. Secondly, we perform RMSNorm before the input of both the attention and feed-forward modules. Thirdly, we use the LLaMA FeedForward module. Additionally, we use Rotary Position Embedding (Su et al., 2021). For context extension, we use Focused Transformer (Tworkowski et al., 2023). Table 7 presents the details about both standard and long-context pretraining.
@@
Parameter/Model Size 270M 3B
Vocab Size 32000 32000
Embedding Size 1024 3200
Num Attention Layers 12 26
Num Attention Heads 8 32
Head Size 128 100
MLP Hidden Size 4096 8640
FoT Context Extension Layers [6, 9] [6, 12, 18]
Table 7: Training details. We pretrain a custom $270$M parameter model and take a pretrained $3$B parameter OpenLLaMAv2 model (Geng, 2023).
Stage Parameter/Model Size 270M 3B
Pretraining Context 2K 2K
Tokens 6.3B 1T
Long Context Pretraining Context 32K 32K
Batch Size 128 128
Start Learning Rate 5e-5 1.5e-5
End Learning Rate 5e-6 1.5e-6
Warmup Steps 250 1000
Tokens 1B 5.4B
Appendix B Detailed results
In this section, we present detailed results on all considered datasets. The training methodology is described in Section 3.1. Details about the number of tokens used for evaluation can be found in Table 14.
Tables 8 and 9 show the results of training $270$M parameter model for $32$K context on a 50/50 mixture of RedPajama data (organized in a standard way) and code data organized using a specified method. Table 8 contains detailed results from training on C# and Python. Table 9 contains results on C averaged across three different subsets of C (for details about the construction of those subsets see Appendix F). Both tables show that SPLiCe outperforms the Baseline by a significant margin and is slightly better than or on par with Repo.
The main advantage of SPLiCe over the Repo approach is that it can be used also for non-structured data. Table 10 shows the detailed results of applying the SPLiCe on non-code data. Note that training on non-code data allows us to improve the model perplexity on the arXiv dataset in comparison to the model trained on code.
Table 11 shows that a simple artificial extension of example length via random concatenation of documents does not help.
Table 8: Perplexity results comparing different ways of organizing the same data. All runs started from the same $270M$ model with $2048$ context and were trained for $32$K context on a 50/50 mixture of RedPajama (organized in a standard way) and code organized in mentioned ways. For details about training please refer to Section 3.1.

### v5

@@ 4 Large-Scale Models
In this section, we show that SPLiCe can improve the long context performance of large-scale language models. To this end, we use $3$B and $7$B parameter models and tasks that test in-context learning (Section 4.2.1), question answering (Section 4.2.2), and in-context information retrieval capabilities (Section 4.2.3). We provide perplexity results in Appendix E.
We also emphasize that those improvements occur during a relatively short, compared to pre-training, fine-tuning. To be more precise, $3$B models were tuned on $5.4$B tokens, whereas $7$B models were tuned on $2$B tokens. All of the models were pre-trained on $1$T tokens.
@@ 4.1 Experimental Setup
For $3$B experiments, we continue training OpenLLaMA $3$B v2 (Geng & Liu, 2023; Geng, 2023) on a 50/50 mixture of RedPajama prepared in the standard way and C prepared using SPLiCeBM25. For $7$B ones, we continue the training of OpenLLaMA $7$B v2 on 75/25/25 mixture of RedPajama (50) prepared in the standard way, StackExchange (25) and C (25) prepared using SPLiCeBM25. StackExchange is part of RedPajama (TogetherComputer, 2023), and C data come from StarCoder (Li et al., 2023b).
We train with $32$K context length using $5.4$B tokens for $3$B models and $2$B for $7$B models (recall that OpenLLaMA models have the $2$K context length). We use a batch size of $256$K tokens per step and the learning rate of ${1.5}\mathrm{e}{-5}$ with linear warmup and cosine decay, following (Geng & Liu, 2023).
Regarding the context extensions method we use FoT (Tworkowski et al., 2023), and the CodeLlama (CL) (Rozière et al., 2023) approach. More explicitly, we test six models:
[…]
where Baseline denotes the standard data preparation method. Hyperparameter details are located in Appendix A.
4.2 Experimental Results
We present perplexity measurements in Appendix E. In particular, following (Anthropic, 2023), we examine perplexity improvements in the function of the token position in the document. Our findings reveal that SPLiCe models demonstrate superior perplexity across all distances. Notably, the improvement in perplexity is more pronounced for tokens positioned later in the document, indicating enhanced language modeling capabilities.
For the remainder of this section, we focus on testing the in-context learning, question answering and information retrieval abilities on our models using downstream tasks.
4.2.1 In-Context Learning
To evaluate the improvements of in-context learning abilities, we use TREC (Li & Roth, 2002; Hovy et al., 2001) and DBpedia (Lehmann et al., 2015), which are text classification tasks. For TREC, to test the long context utilization, we vary the number of in-context examples in $\{190,380,780,1560\}$ (which correspond roughly to $\{4,8,16,32\}$K context length). We sample these examples multiple times and measure classification accuracy on test set. In Table 4, we observe that SPLiCe improves the average performance across different context lengths and model sizes. We follow a very similar protocol for DBpedia, see Table 5, confirming that SPLiCe improves the context utilization. In Appendix I, we study the distribution of improvements with respect to the choice of the in-context examples, showing the domination of SPLiCe.
In Table 4 and Table 5, we report mean improvement $\Delta$ along with $95\%$ bootstrap confidence intervals. See Appendix I for related histograms.
Table 4: Average classification performance on TREC (Li & Roth, 2002; Hovy et al., 2001). For TREC, we average across $50$ sets of in-context examples for $3$B models and $10$ sets for $7$B models. We use $\Delta$ [confidence interval] to denote mean improvement and its $95\%$ bootstrap confidence intervals (see Appendix I).
TREC
Model
Context
(#examples)
Baseline SPLiCe $\Delta$ [conf interv]
[…]
$16$K (780) 68.9 76.9 8.0 [6.9,9.3]{6.9,9.3]}6.9 , 9.3 ]
$8$K (380) 66.5 75.8 9.3 [8.1,10.8]{8.1,10.8]}8.1 , 10.8 ]
$4$K (190) 65.1 72.8 7.7 [6.5,9.0]{6.5,9.0]}6.5 , 9.0 ]
[…]
$16$K (780) 74.0 79.0 5.0 [3.4,6.0]{3.4,6.0]}3.4 , 6.0 ]
$8$K (380) 72.1 76.2 4.1 [2.9,5.7]{2.9,5.7]}2.9 , 5.7 ]
$4$K (190) 66.9 69.7 2.8 [1.5,4.2]{1.5,4.2]}1.5 , 4.2 ]
[…]
$16$K (780) 81.4 82.5 1.1 [0.2,1.6]{0.2,1.6]}0.2 , 1.6 ]
$8$K (380) 76.6 76.5 -0.1 [1.0,1.1]{-1.0,1.1]}- 1.0 , 1.1 ]
$4$K (190) 68.1 69.1 1.0 [0.4,2.9]{-0.4,2.9]}- 0.4 , 2.9 ]
Table 5: Average performance on DBpedia (Lehmann et al., 2015). We average results across $40$ sets of in-context examples, with $\Delta$ [confidence interval] denoting mean improvement and its $95\%$ bootstrap confidence interval. Due to the size of the evaluation dataset for each set of in-context examples, we subsample a subset of $500$ elements of the evaluation set.
DBpedia
Model
Context
(#examples)
Baseline SPLiCe $\Delta$ [conf interv]
[…]
$16$K (190) 79.1 82.0 2.9 [2.5,3.4]{2.5,3.4]}2.5 , 3.4 ]
[…]
$16$K (190) 83.6 85.6 2.0 [1.5,2.5]{1.5,2.5]}1.5 , 2.5 ]
[…]
$16$K (190) 96.2 96.4 0.2 [0.0,0.3]{0.0,0.3]}0.0 , 0.3 ]
Table 6: Average performance on HotPotQA (Yang et al., 2018). We average results across $7$ sets of in-context examples, with with $\Delta$[confidence interval] denoting mean improvement and its $95\%$ bootstrap confidence intervals.
HotPotQA
Model
Context
(#examples)
Baseline SPLiCe $\Delta$ [conf interv]
[…]
4.2.2 Question Answering
We also show that SPLiCe models have improved question answering capabilities. To this end, we use Qasper (Dasigi et al., 2021; Shaham et al., 2022) and HotPotQA (Yang et al., 2018). Qasper contains questions regarding research papers provided as a part of the input context. In Table 7, we observe that SPLiCe improves the two-shot performance. Similarly, we observe improvements on HotPotQA, see Table 6. We stress that in both cases, the results measure in-context capabilities as neither of the datasets was used during training.
@@ 4.2.2 Question Answering
Table 7: Two-shot F1 performance on the validation subset of Qasper (Dasigi et al., 2021). For Qasper, we use the implementation from Language Model Evaluation Harness (Gao et al., 2021). Note that in the $3$B model case, despite using SPLiCe for code data only, we still have improvements in non-code tasks.
QASPER
Model
Context length
(#examples)
Baseline SPLiCe
[…]
4.2.3 Key Retrieval Performance
@@
The key retrieval task introduced in (Liu et al., 2023) has become a routine diagnostic tool to study context capabilities. Specifically, the model is presented with a list of key-value pairs, which are $128$-bit UUIDs, and is tasked to retrieve the value for a given key. Even though very simple, the task proved to be challenging for many open-source language models, see Appendix D for the task details.
In Figure 2, we present the key retrieval performance of $7$B models trained with CodeLlama (Rozière et al., 2023) context extension method. We note that despite relatively short tuning, SPLiCe significantly helps on hard-to-retrieve positions. We provide the results for FoT models in Appendix D.
@@
Figure 2: Key-value retrieval performance on a dictionary of $300$ key-value pairs ($\approx$ $24$K tokens). The $7$B CL model trained with SPLiCe achieves much higher accuracy on hard-to-retrieve positions. The details about this task can be found in Appendix D. Each position was evaluated using $500$ examples.
4.2.4 Shuffling
In our experiments with large $7$B CodeLlama models, we have noted slight improvements in handling long inputs when shuffling documents inside the context, that is, setting Order in the Algorithm 1 to be a random shuffle. We provide a comparison on TREC in Table 17.
@@ 5 Related Work
There is an increasing number of works aiming to study the role of data in LLM training in detail. For instance, (Levine et al., 2022) developed a theory and demonstrated empirically that incorporating non-adjacent but semantically related sentences in training examples leads to better sentence embeddings and improves open-domain question-answering performance. Another work (Gu et al., 2023) introduced a pretraining framework grounded on the idea that text documents often include intrinsic tasks. They showed that this approach substantially boosts in-context learning. Additionally, there is existing work on training long-context language models using repository-level code data, such as (Wu et al., 2022). A recent work (Chan et al., 2022) identifies the training data’s distributional properties that affect transformer models’ in-context capabilities. Similarly, (Han et al., 2023) constructs small-scale data using an iterative gradient approach and shows that such data improve in-context performance.
Our methodology diverges from these works in several key ways. Firstly, we focus on the document-level context during the training phase, as opposed to sentence-level (Levine et al., 2022) or paragraph-level (Gu et al., 2023) granularity. We demonstrate the efficacy of this approach in large-scale language modeling, specifically with OpenLLaMA $3$B. Secondly, we construct a tree structure of related documents through BM25/Contriever-MSMARCO retrieval and linearize this structure to form long-context examples. This allows for greater control over example coherence compared to relying solely on natural data structures like repository-level code. The gradient-based method presented in (Han et al., 2023) can be broadly associated with retrieval used in our work; however, it differs in scale and granularity.
Concurrently to our research, (Shi et al., 2023b) showed a similar method for preparing the training data. However, in our work, we focus on tuning the models for long context, up to $32$K tokens, whereas (Shi et al., 2023b) focuses on training the models from scratch, with context length up to $8$K tokens. In addition to that, we provide ablations regarding context extension methods.
@@ 6 Limitations and Future Work
We show that structuring the training data is a viable way of improving the model’s performance. The presented method can viewed as a general framework for organizing the documents into training examples. This opens multiple further research avenues.
@@
Degree of relevance In this work, SPLiCe constructed training examples by concatenating most matching documents (according to BM25/Cont/Repo). However, recent results of (Tworkowski et al., 2023) show that introducing unrelated data to the context may help the model learn better representations. We leave the study of how the choice of retriever (in particular, the ratio of related and unrelated documents) affects performance for future work. Note that as BM25 is a syntactic retriever, it is not guaranteed that the documents returned by it are semantically related. Moreover, we have not tuned Contriever-MSMARCO, and it was provided only with a $512$ token long prefix of a document.
@@
In order to prepare the training data, SPLiCe uses each document exactly once (we mask out each used document for further retrieval). However, it is possible that allowing some high-quality documents to occur more than once may be beneficial.
Retrieval granularity Another avenue for future work is to study the granularity of the pieces from which the training examples are constructed. In this work, we focus on the document-level granularity. However, it is possible to construct training examples from smaller pieces, such as paragraphs or sentences.
Scaling Our method requires further studies to understand its scaling properties beyond the range presented in the paper.
Other context extension methods In most cases, our models were trained for a long context using the methodology presented in (Tworkowski et al., 2023). Additionally, we tested three popular context extension methods on a medium scale (Naive, YaRN, and CodeLlama) and tuned a large $7$B model using the CodeLlama approach, preliminary confirming the applicability of SPLiCe. However, we leave more detailed studies of the design choices of SPLiCe with other context extension methods to future work.
Other data sources One of the approaches to training long-context language models is to use conversational data (Li et al., 2023a). This is complementary to our method. SPLiCe can utilize data that already exists in vast quantities and can be easily applied to different types of text (like code, Wikipedia articles, StackExchange questions, answers, etc.) to further increase the number of long-context examples. We leave researching how SPLiCe integrates with other methods for preparing the long-context data as future work.
@@
Data curation Using highly correlated samples has the potential to result in training instability. However, we noted no performance degradation during our experiments. We leave the study of how SPLiCe integrates with different data types for the future.
@@ 7 Conclusions
In this work, we present SPLiCe, a novel method of constructing training examples for long-context language models, which utilizes BM25/Contriever-MSMARCO to find relevant documents and feed them to the model in a structured manner. We show that SPLiCe improves the perplexity of the language modeling task and performance on downstream tasks. We further show that SPLiCe can be used to improve long-context utilization of large-scale models using only short fine-tuning. We believe that our work indicates multiple interesting research directions for improving the performance of long-context language models with structured data.
8 Impact Statement
This paper presents work whose goal is to improve the long context utilization ability of language models through structuring training data. It bears standard risks associated with advancing the field of LLMs (such as misuse and generation of malicious/misleading content). For a detailed overview, we refer to (Bender et al., 2021; Weidinger et al., 2021).
Acknowledgments
We gratefully acknowledge that our research was supported with Cloud TPUs from Google’s TPU Research Cloud (TRC). We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Centers: ACK Cyfronet AGH, CI TASK) for providing computer facilities and support within computational grant numbers PLG/2023/016265. Piotr Miłoś was supported by National Science Center Poland under the grant agreement 2019/35/O/ST6/03464.
@@ Appendix A Architecture
The architecture of our models is based on LLaMA (Touvron et al., 2023), and the architectural details can be found in Table 8. Briefly speaking, our architecture is similar to the one introduced in (Vaswani et al., 2017) with a few standard changes. First, we use only the decoder without the encoder part. Secondly, we perform RMSNorm before the input of both the attention and feed-forward modules. Thirdly, we use the LLaMA FeedForward module. Additionally, we use Rotary Position Embedding (Su et al., 2021). For context extension, we use Focused Transformer (FoT)(Tworkowski et al., 2023), CodeLlama (Rozière et al., 2023) (CL) and YaRN (Peng et al., 2023). Table 9 presents the details about both standard and long-context pretraining.
@@
Parameter/Model Size 270M 3B 7B
Vocab Size 32000 32000 32000
Embedding Size 1024 3200 4096
Num Attention Layers 12 26 32
Num Attention Heads 8 32 32
Head Size 128 100 128
MLP Hidden Size 4096 8640 11008
FoT Context Extension Layers [6, 9] [6, 12, 18] [8, 16, 24]
Table 9: Training details. We pretrain a custom $270$M parameter model and take a pretrained $3$B/$7$B parameter OpenLLaMAv2 model (Geng, 2023). Subscript denotes that parameter was specific for a context extension method with FoT referring to (Tworkowski et al., 2023) and no-FoT to other methods (Naive, YaRN (Peng et al., 2023), CodeLlama (Rozière et al., 2023)).
Stage Parameter/Model Size 270M 3B 7B
Pretraining Context 2K 2K 2K
Tokens 6.3B 1T 1T
Long Context Pretraining Context 32KFoT, 16Kno-FoT 32K 32K
Batch Size 128FoT, 16no-FoT 128 128FoT, 32no-FoT
Start Learning Rate 5e-5 1.5e-5 1.5e-5
End Learning Rate 5e-6 1.5e-6 1.5e-6
Warmup Steps 250 1000 1000FoT, 200no-FoT
Tokens 1B 5.4B 2B
Appendix B Detailed Results For Medium-Size Models
In this section, we extend results presented in Section 3. The training methodology is described in Section 3.1. Details about the number of tokens used for evaluation can be found in Table 18.
Tables 10 and 11 show the results of training $270$M parameter model for $32$K context on a 50/50 mixture of RedPajama data (organized in a standard way) and code data organized using a specified method. Table 10 contains detailed results from training on C# and Python. Table 11 contains results on C averaged across three different subsets of C (for details about the construction of those subsets, see Appendix G). Both tables show that SPLiCe outperforms the Baseline by a significant margin.
The main advantage of SPLiCeBM25/SPLiCeCont over the SPLiCeRepo approach is that it can be used also for non-structured data. Table 12 shows the detailed results of applying the SPLiCe on non-code data. Note that training on non-code data allows us to improve the model perplexity on the arXiv dataset in comparison to the model trained on code.
Table 13 considers models trained with YaRN (Peng et al., 2023), CodeLlama (Rozière et al., 2023) and Naive (no adjustment to RoPE) context extension methods. Table 14 shows that a simple artificial extension of example length via random concatenation of documents does not help.
Table 10: Perplexity results comparing different ways of organizing the same data. All runs started from the same $270M$ model with $2048$ context and were trained for $32$K context on a 50/50 mixture of RedPajama (organized in a standard way) and code organized in the mentioned ways. For details about training please refer to Section 3.1.